OpenStudy (faolan):

2x - 4y = 10 7x - 9y = 12 HOW DO I SOLVE THIS?

OpenStudy (anonymous):

multiply the top equation by 7 and the bottom equation by 2, so 2x and 7x are then both 14x. Then you can cross them out, solve for y, then with y, pick one of the equations (dosent matter which) and solve for x.

OpenStudy (faolan):

Nvm i got it!! x= -23/5 and y= 21/5

OpenStudy (anonymous):

nice